OpenStudy (iheart.jesus):

The table shows a function. Is the function linear or nonlinear? x y 7 16 8 11 9 6 linear or nonlinear

OpenStudy (herb12):

That function is linear because it have exactly one output for each input.

OpenStudy (campbell_st):

its linear since the change in y values is constant...

OpenStudy (kyraarama):

Yes, the function is linear and I'll explain why. The general formula for a linear function is: f(x)= mx +t So we get the first two points and apply them to the formula: 1. 7m + t = 16 2. 8m + t = 11 So we subtract the second one of the first one: 1-2. 16- 11 = 8m + t - 7m -t -> 5 = -m -> -5 = m in 1. 7 * (-5) + t = 16 -> t = 51 So the formula is: f(x) = -5x + 51 Now we controll this with the last point: -5 * 9 + 51 = 6 The formula is correct and the table gives us three points of a linear function. I hope I could help you and I am sorry for my english (not my mother tongue).